OpenStudy (anonymous):

1. A doctor needs to prescribe a certain medicine for her patient. The amount of medicine prescribed per day is based on the weight of the patient at a ratio of three milligrams per two pounds of bodyweight. If the patient weighs 150 pounds, how much medication should the doctor prescribe per day?

OpenStudy (anonymous):

set up a proportion, defining a variable for the unknown quantity, b) solve the proportion, showing all work leading to your answer, and c) write a sentence to explain the meaning of the answer for the situation thats the directions

OpenStudy (imstuck):

set it up as a ratio. Like this:\[\frac{ mg med }{lb }=\frac{ mg med }{ lb }\]fill in your ratios like this:\[\frac{ 3mg }{ 2lb }=\frac{ x }{ 150lb }\]

OpenStudy (imstuck):

Cross multiply to find x, which is the amount of medication for a person weighing 150 lbs.

OpenStudy (imstuck):

3(150) = 2x 450=2x x=???

OpenStudy (anonymous):

225

OpenStudy (imstuck):

225 mg of medicine is correct.

OpenStudy (anonymous):

yay thanks